OpenStudy (study_buddy99):

Desperate to learn Alg 2:Simplify (4r^0)^4

OpenStudy (solomonzelman):

In general, when \(\color{#000000}{ \displaystyle x\ne0 }\), then \(\color{#000000}{ \displaystyle x^0=1 }\).

OpenStudy (study_buddy99):

so 4^4?

OpenStudy (solomonzelman):

I will give you a similar example: \(\color{#000000}{ \displaystyle (5s^0)^3 =(5\cdot 1)^3=5^3=125 }\)

OpenStudy (solomonzelman):

Yes, exactly!

OpenStudy (study_buddy99):

amazing, Thank you so much!

OpenStudy (solomonzelman):

Not a problem, you have done it yourself:)